OpenStudy (anonymous):

if f(x)=2x+5 and f(a)=-1 then what is a

OpenStudy (anonymous):

f(x) = 2x + 5 and f(a) = -1 substitute a into f(x) f(a) = 2a + 5 Substitue f(a) = -1 -1 = 2a + 5 -6 = 2a a = -3.

OpenStudy (anonymous):

yes it is a right answer me=yeung .

OpenStudy (anonymous):

thank you thank you